# DPoser: Diffusion Model as Robust 3D Human Pose Prior

### 09 Dec 2023 v1

@@ 4.2 Optimization Leveraging Diffusion Priors
where $M_{J}(\theta,\beta)$ and $\Pi_{C}$ symbolize the 3D joint positions regressed from SMPL and the 3D-to-2D camera projection function, respectively. $J^{\text{est}}$ are the 2D keypoints estimated using an off-the-shelf 2D pose estimator (OpenPose [60] in our experiments), and $\lambda_{i}$ indicates the confidence score of joint $i$. The Geman-McClure error function [61] $\rho$ serves to robustly quantify discrepancies in 2D joint positions.
The bending term $L_{\theta}$ is incorporated to penalize excessive bending at the elbows and knees, formulated as $L_{\theta}=\sum_{i\in\text{(elbows, knees)}}\exp(\boldsymbol{\theta}_{i})$. The shape regularizer $L_{\beta}$ is a simple L2 norm, i.e., $L_{\beta}=\|\beta\|_{2}^{2}$. The constant weights for prior terms are denoted as $w_{\theta},w_{\beta}$ and $w_{\alpha}$, respectively.
Contrasting SMPLify’s GMM-based prior, we employ $L_{\mathrm{DPoser}}$ to draw the pose closer to our learned manifold. Given the structure of $L_{\mathrm{DPoser}}$ (as seen in Eq. (8)), a crucial aspect lies in judiciously selecting the diffusion timestep $t$ during the iterative optimization process. We address this concern by introducing our novel truncated timestep scheduling strategy in the subsequent section.
@@ 4.3 Test-time Truncated Timestep Scheduling
1:A trained diffusion model $\epsilon_{\theta}(\mathbf{x}_{t};t)$, task-specific loss $L_{\text{task}}$, range of diffusion timesteps $[t_{\text{max}},t_{\text{min}}]$, number of optimization iterations $N$, DPoser weights $w_{t}$
2:Initialization of parameters $\mathbf{x}_{0}$
@@
7:     $\mathbf{\hat{x}}_{0}(t)\leftarrow\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\theta}(\mathbf{x}_{t};t)}{\alpha_{t}}$triangleright One-step denoiser
8:     $L_{\text{DPoser}}\leftarrow w_{t}\lVert\mathbf{x}_{0}-\text{sg}[\mathbf{\hat{x}}_{0}(t)]\rVert_{2}^{2}$
@@
Insights from pose generation. In optimizing pose-related tasks, one fundamental challenge arises from the gap between the image-based data and the articulated pose parameters we manipulate. Prior studies in image domains [15] have noted that earlier timesteps (i.e., smaller $t$) are typically aligned with perceptual content, while the later timesteps are more closely associated with the refinement of details in the diffusion process. As illustrated in Fig. 3, our experimental foray into employing diffusion models for pose generation yields an interesting insight into pose domains: the concentration of meaningful information resides at the “clean” side of the diffusion trajectory, rather than being dispersed uniformly from noise to clean states. This observation serves as the catalyst for our introduction of a truncated annealed timestep scheduling mechanism.
Truncated timestep scheduling. Departing from traditional diffusion models that uniformly sample across the continuum from $1.0$ to $0.0$, we introduce a selective sampling strategy. Specifically, we truncate the timestep range to intervals such as $0.2\to 0$, thereby focusing on the data segments most laden with meaningful information. Based on the linear annealed scheduling, the truncated timestep $t$ for each optimization step can be formulated as:
Here, $N$ signifies the total optimization iterations, while iter represents the current optimization iteration. We summarize our DPoser optimization framework as Algorithm 1, where the truncated range $[t_{\text{max}},t_{\text{min}}]$ is typically set to $[0.2,0.001]$.
Through ablation studies, we prove that the proposed truncated timestep scheduling offers significant advantages over the random, fixed, and full annealed ones utilized in previous works [62, 18, 19, 8], delivering more efficient optimization across a wide range of downstream tasks. This is achieved by concentrating computation on the most informative portions of the data trajectory, as informed by our preliminary sampling investigations on pose generation.
@@ 5 Experiments
In this section, we first elaborate on the implementation details and evaluation metrics. We showcase the robustness and versatility of DPoser across a spectrum of pose-centric applications, including pose generation, human mesh recovery, pose completion, and motion denoising. Additionally, we conduct an ablation study to delve into the key elements contributing to our model’s performance.
@@
Table 1: Comparative analysis of pose generation metrics. The discrepancy between visual impressions and APD/SI metrics is discussed, with reference to Fig. 4. *Indicates the use of a DDIM sampler [11] with a reduced 10-step discretization.
(a) AMASS samples
(b) DPoser (ours)
(c) DPoser (ours)*
(d) GMM
(e) Pose-NDF
(f) VPoser
Figure 4: Qualitative comparison of generated human poses: (b) illustrates naturalistic poses closely aligned with real-world data, whereas (c) shows poses that, despite superior APD and SI metrics, lack natural appearance. *We use a DDIM sampler [11] with only 10 steps.
5.1 Experimental Setup
Implementation details. We train our DPoser model on the AMASS dataset [52], adhering to the same training partition as previous works [35, 36]. The model employs axis-angle representation for joint rotations, which we normalize to have zero mean and unit variance. The architecture consists of a fully connected neural network with approximately 8.28M parameters. It draws inspiration from GFPose [43] but omits conditional input pathways for our unconditional setting. To stabilize training, we use an exponential moving average with a decay factor of 0.9999, as advised by ScoreSDE [8]. The Adam optimizer, a learning rate of $2\times 10^{-4}$, and a batch size of 1280 govern the optimization process. The training of 800,000 iterations takes roughly 8 hours on a single Nvidia RTX 3090Ti GPU.
Evaluation metrics. We adopt task-specific metrics:
  • Pose Generation: Diversity and fidelity are evaluated using Average Pairwise Distance (APD) [37] and Self-Intersection rates (SI) [36], respectively.
  • Human Mesh Recovery: Metrics include Mean Per Joint Position Error (MPJPE) and Procrustes-aligned MPJPE (PA-MPJPE).
  • Pose Completion: We utilize both MPJPE and Mean Per-Vertex Position Error (MPVPE), for masked areas.
  • Motion Denoising: Metrics are MPVPE and MPJPE.
All errors are reported in millimeter units.
5.2 Pose Generation
@@
Interestingly, quantitative metrics such as APD and SI (Tab. 1) do not always corroborate our qualitative findings. For instance, a 10-step DDIM sampler [11]—suboptimal by design—outperformed real-world data [52] in APD, which we attribute to the generation of exaggerated poses.
In summary, our findings underscore the need for a balanced evaluation strategy that merges quantitative metrics with qualitative observations. Beyond generation, DPoser also performs well in interpolation tasks, based on corresponding ODEs with marginal densities akin to Eq. (2[8]. Additional experiments are detailed in the Appendix.
Methods PA-MPJPE $\downarrow$ MPJPE $\downarrow$
GMM [34] 57.53 79.66
VPoser [35] 61.10 85.25
Pose-NDF [36] 61.67 85.09
DPoser (ours) 57.28 84.79
CLIFF [51] 57.30 92.86
+ GMM [34] 55.54 78.32
+ VPoser [35] 56.12 79.68
+ Pose-NDF [36] 55.50 78.39
+ DPoser (ours) 54.87 77.62
Table 2: Performance comparison of human mesh recovery on the EHF dataset [35] using two initialization methods.
Methods Occ. left leg Occ. legs Occ. arms Occ. trunk
MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
Pose-NDF [36] ($S=1$) 499.35 467.26 503.66 470.82 378.61 312.16 319.65 208.35
Pose-NDF [36] ($S=1$${}^{\dagger}$ 597.72 558.45 606.68 566.01 370.3 283.43 346.18 228.54
VPoser [35] ($S=1$) 465.98 437.92 464.38 436.06 387.14 293.17 298.63 195.16
VPoser [35] ($S=1$${}^{\dagger}$ 194.44 177.01 199.78 183.34 205.17 160.60 75.42 48.30
VPoser [35] ($S=5$) 458.92 431.31 456.01 427.90 374.32 283.90 288.96 188.26
VPoser [35] ($S=10$) 455.44 427.79 450.88 423.04 371.26 281.62 286.55 186.28
DPoser (ours) ($S=1$) 85.46 80.66 107.07 100.97 108.27 84.37 46.37 29.99
DPoser (ours) ($S=5$) 48.61 44.79 74.67 69.64 75.92 59.90 26.59 16.68
DPoser (ours) ($S=10$) 39.77 36.15 65.24 60.54 66.77 52.75 21.58 13.41
Table 3: Performance metrics for pose completion under varying occlusion scenarios. $S$ denotes the number of hypotheses. ${}^{\dagger}$ Comparison methods are implemented in the optimization paradigm.
@@ 5.3 Human Mesh Recovery
Figure 5: Qualitative evaluations of HMR leveraging DPoser as pose prior. CLIFF [51] serves as the optimization initializer.
We probe the efficacy of DPoser in HMR, focusing on estimating human pose and shape from monocular images. We conduct experiments on the EHF dataset [35] and benchmark our method against existing SOTA priors. Our optimization-based framework incorporates two initialization paradigms: (1) a baseline initialization that utilizes mean pose values and a default camera setup, and (2) an advanced initialization scheme that leverages CLIFF [51], a pre-trained regression-based model tailored for HMR.
As shown in Tab. 2, DPoser demonstrates robust performance in HMR. With baseline initialization, it closely rivals the strong constraints of the GMM prior [34]. When enhanced by advanced initialization via CLIFF, DPoser surpasses established deep priors like Pose-NDF [36]. Fig. 5 visually confirms DPoser’s superior efficacy and adaptability across multiple datasets including EHF [35], MSCOCO [54], 3DPW [55], and UBody [53]. Further qualitative comparisons are available in the Appendix.
Methods AMASS [52] HPS [56]
VPoser [35] 23.42 22.78
Pose-NDF [36] 22.13 23.82
MVAE [40] 26.80 N/A
HuMoR [39] 22.69 N/A
DPoser (ours) 19.87 20.54
Table 4: Performance metrics (MPJPE) for motion denoising.
@@ 5.4 Pose Completion
Figure 6: Qualitative evaluations of pose completion. Three hypotheses are drawn from VPoser [35] and our DPoser.
@@
PA-MPJPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
Random 68.21 98.57 109.53 102.37 43.33 23.87
Fixed 57.72 85.11 40.44 36.70 45.69 22.54
Full annealed 61.15 85.27 42.97 39.42 39.72 20.80
Truncated annealed 57.28 84.79 39.77 36.15 38.21 19.87
Table 5: Comparative performance of different timestep scheduling across pose-related tasks.
Noise std AMASS [52] HPS [56]
20.00 31.93/13.64 31.93/13.45
40.00 63.81/19.87 63.81/20.54
100.00 159.78/33.18 159.78/35.32
Table 6: DPoser in motion denoising under varying noise scales. MPJPE is reported as before/after applying DPoser denoising.
In practical scenarios like those encountered in the UBody dataset [53] (refer to Fig. 4(d)), HMR algorithms often grapple with occlusions leading to incomplete 3D pose estimates. In this context, our ambition is to recover full 3D poses from such partially observed data, where the unseen parts are initialized with random noise.
Given that Pose-NDF [36] and VPoser [35] have not been previously scrutinized for this task, we introduce two experimental paradigms—generation-based and optimization-based. In the optimization regimen, we incorporate an L2-norm fidelity term for the observed pose and constrain the entire pose using these deep priors. In contrast, the generation-based scheme for VPoser involves encoding the whole pose, followed by decoding and sampling from the resulting Gaussian distributions. For Pose-NDF, we iteratively project the whole pose onto the learned manifold, replacing the unmasked components with known data. For DPoser, we report optimization-based outcomes while the generation-based results employing other inverse problem solvers [8, 20, 19] can be checked in the Appendix.
Given the task’s inherent ambiguity, we sample multiple solutions and select the best-fitting one, except the deterministic Pose-NDF. Tab. 3 and Fig. 6 demonstrate DPoser’s superiority in pose completion under occlusions, offering multiple plausible solutions, unlike VPoser [35] and Pose-NDF [36], with the latter facing generalization issues when applied to such unseen noisy poses.
5.5 Motion Denoising
Though not initially designed for temporal tasks, DPoser shows remarkable proficiency in motion denoising. The task aims to estimate clean body poses from noisy 3D joint positions in motion capture sequences. Adhering to the setup outlined in HuMoR [39], we utilize 60-frame sequences from the AMASS [52] test dataset and artificially introduce Gaussian noise with a standard deviation of 40 mm to the 3D joint positions. Moreover, we conduct experiments on HPS datasets [56] without additional training to validate the generalization. Our optimization loss is composed of a data fidelity term, a temporal consistency term, and our pose prior, as discussed in the Appendix.
As presented in Tab. 4, DPoser sets a new standard in motion denoising, outperforming even specialized motion priors like HuMoR [39]. To further confirm the robustness of DPoser, we conduct evaluations under varying conditions to gauge DPoser’s denoising capabilities. The results, detailed in Tab. 6, reveal that DPoser consistently achieves significant reductions in MPJPE, maintaining robust performance even under extreme noise conditions, specifically at a noise standard deviation of 100.00 mm. We direct the reader to our Appendix for qualitative assessments.
5.6 Ablation Study
In our ablation study, we initially focus on the impact of truncated timestep scheduling on DPoser’s performance. This involves contrasting our proposed scheduling strategy against three established methods—random, fixed, and full annealed scheduling [62, 18, 19, 8]. As Tab. 5 demonstrates, our strategy consistently outperforms these alternatives across all evaluated tasks. Further, we delve into the training aspects of DPoser, such as rotation representations and the integration of an auxiliary loss akin to HuMoR [39]. Additionally, using the same trained prior, we benchmark DPoser against SOTA diffusion-based inverse problem solvers [8, 20, 19] on the pose completion task. The results underscore that DPoser not only proves versatile across multiple pose-related tasks but also excels in performance. Detailed findings and analyses from these ablation studies are presented in the Appendix.

### 23 Mar 2024 v2

@@ 3.3 Optimization Leveraging Diffusion Priors
where $M_{J}(\theta,\beta)$ calculates the 3D joint coordinates through SMPL’s forward kinematics. The function $\Pi_{C}$ maps these 3D coordinates into 2D space, aligning with the camera’s perspective. $J^{\text{est}}$ refers to the 2D keypoints estimated using an off-the-shelf 2D pose estimator (in our case, ViTPose [55]), with $\lambda_{i}$ reflecting the confidence score for each joint $i$. The Geman-McClure error function ($\rho$) is employed to assess the discrepancy in 2D joint locations reliably.
To mitigate the issue of overfitting, which often leads to unrealistic poses when solely minimizing reprojection loss, several regularization terms are introduced. Specifically, alongside our body prior $L_{\mathrm{DPoser}}$, the bending term $L_{\theta}$ is incorporated to penalize excessive bending at the elbows and knees, formulated as $L_{\theta}=\sum_{i\in\text{(elbows, knees)}}\exp(\boldsymbol{\theta}_{i})$. Additionally, the shape regularization term $L_{\beta}=\|\beta\|_{2}^{2}$ is employed to maintain the body shape within plausible bounds. The weights for prior terms are denoted as $w_{\theta},w_{\beta}$ and $w_{\alpha}$, respectively.
Given the structure of $L_{\mathrm{DPoser}}$ (as seen in Eq. (8)), a crucial aspect lies in judiciously selecting the diffusion timestep $t$ during the iterative optimization process. In the subsequent section, we address this concern by introducing our novel truncated timestep scheduling strategy.
@@ 3.4 Test-time Truncated Timestep Scheduling
Motivation from pose generation. Adapting techniques from the image domain to pose data requires a nuanced understanding of the differences between the two. Previous image-based research [4] shows that initial timesteps (larger $t$) correspond to the perceptual content, while later timesteps refine details. Pose data, however, lacks this structured layering and spatial redundancy, indicating a need for a tailored timestep approach in the diffusion process.
@@
As depicted in Fig. 3, we find that pose generation doesn’t benefit from the early timesteps as image generation does. The significant stages of pose refinement occur at smaller $t$, specifically when $t\leq 0.3$. A uniform distribution of timesteps, as tested in (b) with only five steps, proves less effective for pose data. In contrast, allocating these steps toward the latter end of the diffusion process, as in (c), yields significantly better samples, implying the critical information is not evenly distributed but rather is concentrated toward the end.
Truncated timestep scheduling. Based on these insights, we propose a shift from standard uniform timestep sampling to a truncated strategy, especially for pose data. By focusing on the last timesteps, particularly between 0.2 and 0.0, we target the interval rich in pose-specific information. Specifically, based on the linear descending scheduling, the truncated timestep $t$ for each optimization step can be expressed as:
where $N$ denotes the total number of optimization iterations, and iter signifies the current iteration. This formulation is integral to our proposed optimization framework, which is comprehensively summarized in Algorithm 1. The practical implementation typically involves setting the truncated range to $[0.2,0.05]$.
@@
1:A trained diffusion model $\epsilon_{\phi}(\mathbf{x}_{t};t)$, task-specific loss $L_{\text{task}}$, range of diffusion timesteps $[t_{\text{max}},t_{\text{min}}]$, number of optimization iterations $N$.
2:Initialization of SMPL body pose parameters $\mathbf{x}_{0}$
@@
7:     $\mathbf{\hat{x}}_{0}(t)\leftarrow\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\phi}(\mathbf{x}_{t};t)}{\alpha_{t}}$triangleright One-step denoiser
8:     $L_{\text{DPoser}}\leftarrow w_{t}\lVert\mathbf{x}_{0}-\text{sg}[\mathbf{\hat{x}}_{0}(t)]\rVert_{2}^{2}$triangleright DPoser regularization
@@ 4 Experiments
In this section, we showcase the robustness and versatility of DPoser across a spectrum of pose-centric tasks, including pose generation, human mesh recovery, pose completion, and motion denoising. Due to the page limit, we leave experimental details and more qualitative assessments in the Appendix.
4.1 Experimental Setup
Implementation details. We train our DPoser model on the AMASS dataset [32], adhering to the same training partition as previous works [37, 49]. The model employs axis-angle representation for joint rotations, which we normalize to have zero mean and unit variance. The architecture consists of a fully connected neural network with approximately 8.28M parameters. It draws inspiration from GFPose [8] but omits conditional input pathways for our unconditional setting. To stabilize training, we use an exponential moving average with a decay factor of 0.9999, as advised by [48]. The Adam optimizer, a learning rate of $2\times 10^{-4}$, and a batch size of 1280 govern the optimization process. The training of 800,000 iterations takes roughly 8 hours on a single Nvidia RTX 3090Ti GPU.
Evaluation metrics. To comprehensively evaluate our models across various tasks, following Pose-NDF [49], we adopt task-specific metrics:
  • Pose Generation: Diversity and fidelity are evaluated using Average Pairwise Distance (APD) and Self-Intersection rates (SI), respectively.
  • Human Mesh Recovery: The Procrustes-aligned Mean Per Joint Position Error (PA-MPJPE) measures the accuracy of recovered human meshes.
  • Pose Completion: The Mean Per Joint Position Error (MPJPE) for masked body joints serves as the metric, focusing on the inferred occluded parts.
  • Motion Denoising: Both MPJPE and the Mean Per-Vertex Position Error (MPVPE) are calculated to assess the denoising effectiveness.
All errors are reported in millimeter units.
4.2 Pose Generation
(a) GAN-S [9]
(b) DPoser (ours)
(c) DPoser (ours)*
(d) GMM [1]
(e) Pose-NDF [49]
(f) VPoser [37]
Figure 4: Qualitative comparison of generated human poses: (b) illustrates naturalistic poses aligned with real-world data, whereas (c) shows poses that, despite superior metrics, lack natural appearance. *We use a DDIM sampler [44] with only 10 steps.
@@ 4.2 Pose Generation
Table 1: Comparative analysis of pose generation metrics. The discrepancy between visual impressions and APD/SI metrics is discussed, with reference to Fig. 4. *Indicates the use of a reduced 10-step sampler.
@@
Interestingly, quantitative metrics such as APD and SI (Tab. 1) do not always corroborate our qualitative findings. For instance, a 10-step DDIM sampler [44]—suboptimal by design—outperformed real-world data [32] in APD, which we attribute to the generation of exaggerated poses. In summary, our findings underscore the need for a balanced evaluation strategy that merges quantitative metrics with qualitative observations.
@@ 4.3 Human Mesh Recovery
Initialization No fitting GMM [1] VPoser [37] Pose-NDF [49] GAN-S [9] DPoser(Ours)
from scratch 108.57 58.32 58.08 57.87 57.26 56.05
CLIFF [25] 56.62 51.02 49.39 49.50 49.58 49.05
Table 2: Performance comparison of human mesh recovery on the EHF dataset [37] using two initialization methods. PA-MPJPE is reported as the metric.
Figure 5: Human mesh recovery. (a) Fitting from scratch. *Ground truth for the EHF dataset is annotated in SMPL-X [37], which extends SMPL [31] with fully articulated hands and an expressive face. (b) Initialization using the CLIFF [25] prediction.
@@
Figure 6: Qualitative evaluations of human mesh recovery leveraging DPoser as pose prior. CLIFF [25] serves as the optimization initializer.
We probe the efficacy of DPoser in human mesh recovery (HMR), focusing on estimating human pose and shape from monocular images. We conduct experiments on the EHF dataset [37] and benchmark our method against existing SOTA priors. Our optimization-based framework incorporates two initialization paradigms: (1) a baseline initialization that utilizes mean pose values and a default camera setup, and (2) an advanced initialization scheme that leverages CLIFF [25], a pre-trained regression-based model tailored for HMR. Moreover, GAN-S [9] implementations require a GAN-inversion phase to convert initial poses into their latent representations, which is notably time-consuming.
Tab. 2 and Fig. 5 showcase the comparative performance of DPoser, highlighting its exceptional ability in HMR tasks. Notably, when fitting from scratch, it surpasses established SOTA priors like GAN-S [9] and Pose-NDF [49] and rivals the specific regression-based model [25]. The integration of CLIFF as initialization further amplifies DPoser’s performance, underscoring its efficiency and the benefits of employing refined starting conditions. Fig. 6 further confirms DPoser’s superior efficacy and adaptability across multiple datasets including EHF [37], MSCOCO [27], 3DPW [51], and UBody [26].
@@ 4.4 Pose Completion
In practical scenarios like those encountered in the UBody dataset [26] (refer to Fig. 5(d)), HMR algorithms often grapple with occlusions leading to incomplete 3D pose estimates. In this context, our ambition is to recover full 3D poses from partially observed data, initializing the occluded parts with random noise. Our DPoser model is employed to refine these initially implausible poses into feasible ones, utilizing an L2 loss on the visible parts to ensure data consistency.
Initialization VPoser Pose-NDF DPoser
Zeros 180.90 157.50 73.92
10mm noise 181.86 172.50 74.69
100mm noise 180.25 511.51 74.19
Table 3: Pose completion on the AMASS [32] dataset (left leg under occlusion, single-hypotheses) using various initialization strategies. DPoser demonstrates its effectiveness across all conditions.
In parallel, we employ a comparable optimization strategy for both Pose-NDF [49] and VPoser [37]. Notably, Tab. 3 reveals that Pose-NDF struggles with poorly initialized poses unseen during its training phase. To mitigate this issue, we have to initialize the occluded poses near zero (close to rest pose) for Pose-NDF to prevent optimization divergence. Additionally, as a task-specific baseline, we adapt the original VPoser model into CVPoser by incorporating conditional inputs within its VAE framework [22]. This modification enables the encoder and decoder to process additional partial poses, facilitating end-to-end conditional sampling.
Methods Occ. left leg Occ. legs Occ. arms Occ. trunk
PoseNDF ($S=1$[49] 158.21 159.19 201.00 75.42
PoseNDF ($S=5$) 147.66/158.11/7.62 151.86/159.21/5.33 196.36/200.92/3.30 70.88/75.39/3.25
PoseNDF ($S=10$) 144.38/158.06/8.31 149.38/159.14/5.90 194.79/200.87/3.63 69.45/75.38/3.54
VPoser (S=1) [37] 180.78 198.18 159.86 37.75
VPoser ($S=5$) 167.92/181.30/10.53 178.77/198.15/14.51 148.17/159.65/8.64 31.83/37.79/4.54
VPoser ($S=10$) 162.82/181.09/12.21 172.83/198.31/16.30 144.53/159.80/9.69 30.06/37.78/4.99
CVPoser ($S=10$) ${}^{\dagger}$ 71.66/145.52/51.68 90.49/148.30/38.46 83.02/136.82/36.47 18.77/37.83/13.12
DPoser(ours) ($S=1$) 74.48 97.39 81.49 28.58
DPoser(ours) ($S=5$) 42.64/73.85/24.36 67.70/97.06/22.29 58.52/82.37/18.33 17.11/28.59/8.92
DPoser(ours) ($S=10$) 35.37/74.01/26.47 59.25/96.77/24.55 51.27/81.76/20.04 13.95/28.57/9.85
Table 4: Performance metrics (min/mean/std of MPJPE across multiple hypotheses) for pose completion under varying occlusion scenarios. $S$ denotes the number of hypotheses. ${}^{\dagger}$ Task-specific baseline trained with partial poses as conditional input.
Figure 7: Qualitative evaluations of pose completion. Three hypotheses are drawn for each method. DPoser uniquely offers multiple plausible solutions for partially observed poses, a scenario where competitors often struggle due to limited generalization.
Given the inherent uncertainties within this task, we generate multiple solutions and evaluate them based on their minimum, mean, and standard deviation errors against the ground truth. As illustrated in Tab. 4, DPoser exhibits superior performance across different occlusion scenarios compared to existing pose priors and even the task-specific CVPoser, highlighting its effectiveness in pose completion. The qualitative evaluations are presented in Fig. 7. Here, we observe that DPoser can generate a multitude of plausible poses, a capability lacking in VPoser [37]. Pose-NDF [49], meanwhile, struggles with generalizing to unseen noisy poses and making plausible adjustments from its rest pose initialization.
4.5 Motion Denoising
Though not initially designed for temporal tasks, DPoser shows remarkable proficiency in motion denoising. The task aims to estimate clean body poses from noisy 3D joint positions in motion capture sequences. Adhering to the setup outlined in HuMoR [40], we utilize 60-frame sequences from the AMASS [32] dataset and artificially introduce Gaussian noise with a standard deviation of 40 mm to the 3D joint positions. Moreover, we conduct experiments on HPS datasets [15] without additional training to validate the generalization.
As presented in Tab. 5, DPoser sets a new standard in motion denoising, outperforming even specialized motion priors like HuMoR [40]. To further confirm the robustness of DPoser, we conduct evaluations under varying conditions to gauge DPoser’s denoising capabilities. The results, detailed in Tab. 6, reveal that DPoser consistently achieves significant reductions in MPJPE, maintaining robust performance under extreme noise conditions.
4.6 Ablation Study
Methods AMASS [32] HPS [15] No prior 24.19 23.67 VPoser [37] 23.42 22.78 Pose-NDF [49] 22.13 21.60 MVAE [28] 26.80 N/A HuMoR [40] 22.69 N/A DPoser (ours) 19.87 20.54 Table 5: Performance metrics (MPJPE) for motion denoising under 40 mm noise.
Noise std AMASS [32] HPS [15] 20.00 31.93/13.64 31.93/13.45 40.00 63.81/19.87 63.81/20.54 100.00 159.78/33.18 159.78/35.32 Table 6: DPoser in motion denoising under varying noise scales. MPJPE is reported as before/after applying DPoser denoising.
@@ 4.6 Ablation Study
PA-MPJPE $\downarrow$ MPJPE ($S=10$) $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
Random 58.84 86.23/121.57/23.16 43.33 23.87
Fixed 56.55 36.99/71.68/23.41 45.69 22.54
Uniform 59.28 42.72/75.70/21.84 39.72 20.80
Truncated 56.05 35.37/74.01/26.47 38.21 19.87
Table 7: Evaluation of timestep scheduling strategies on key pose-related tasks, highlighting the superior efficacy of the proposed truncated scheduling.
In our ablation study, we initially focus on the impact of truncated timestep scheduling on DPoser’s performance. This involves contrasting our proposed scheduling strategy against three established methods—random, fixed, and uniform scheduling [34, 33, 6, 48]. As Tab. 7 demonstrates, our strategy consistently outperforms these alternatives across all evaluated tasks. Additionally, we delve into the training aspects of DPoser, such as rotation representations and the integration of an auxiliary loss akin to HuMoR [40]. Using the same trained prior, we also compare DPoser’s capabilities with SOTA diffusion-based solvers [48, 7, 6] on pose completion, revealing its superior versatility and performance. Detailed findings and analyses from these ablation studies are presented in the Appendix.